A software designer is mapping the streets for a new racing game. All of the streets are depicted as either perpendicular or parallel lines. The equation of the lane passing through A and B is -7x + 3y = -21.5. What is the equation of the central street PQ?


1
Expert's answer
2021-07-01T09:13:18-0400


equition of line passing trough A and B "-7x + 3y = -21.5" can also be written as "y=\frac{7}{3}x-\frac{21.5}{3}"

comparing with standard form "y=mx+c"

slope "m=\frac{7}{3}"

PQ is perpendicular to lane AB. Product of slope of perpendicular lines will be -1.

"m_1.m_2=-1"

slope of the line"=\frac{-1}{slope \space of\space parallel\space line}=\frac{-1}{\frac{7}{3}}=\frac{-3}{7}"

equition of the line will then be

"y=mx+c\\y=-\frac{3}{7}x+c\\7y=-3x+7c\\7y+3x=7c"

dividing by 2

"3.5y+1.5x=3.5c"


find c




the line pq passes through the point (7,6) in as shown in the figure hence line could be found out using point slope form of line.


"y-y_1=m(x-4)\\slope=\frac{-3}{7}\space (7,6)\\y-6=\frac{-3}{7}(x-7)=7\\7(y-6)=-3(x-7)\\7y-42-3x+21\\=7x+3x=21+42\\7x+3x=63" is the eqition of central line.

divide by 2

"3.5y+1.5x=31.5"